OpenStudy (anonymous):

What is the difference between (x + y)^2 and x^2 + y^2

OpenStudy (anonymous):

2xy

OpenStudy (anonymous):

Its like This x=2 and y=3 (x+y)^2 = (2+3)^2 =5^2=25 x^2+y^2=2^2+3^2=4+9=13 hope u get it

OpenStudy (anonymous):

(x+y)^2 = x^2 + 2xy + y^2 (x+y)^2 - (x^2+y^2) = (x^2 + 2xy + y^2) - (x^2 + y^2) = x^2 + 2xy + y^2 - x^2 - y^2 = 2xy